A STUDY OF THE A^{2}Π/B^{2}Σ^{+} - X^{2}Σ^{+} BAND SYSTEMS OF CALCIUM MONOHYDRIDE, CaH, USING A SUPERSONIC MOLECULAR BEAM SOURCE AND LASER INDUCED FLUORESCENCE DETECTION

\underline{JAMIE GENGLER}, TIMOTHY STEIMLE, and JINHAI CHEN, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The high-resolution laser induced fluorescence spectrum of a molecular beam sample of calcium monohydride, CaH, in the region of the strongly overlapped (1,0) A^{2}Π-X^{2}Σ^{+} and (0,0) B^{2}Σ^{+}-X^{2}Σ^{+} band systems near 630 nm and the (0,0) A^{2}Π-X^{2}Σ^{+} band system near 690 nm have been recorded and analyzed. The spectral features exhibit a small splitting which is attributed to proton magnetic hyperfine interactions in the X^{2}Σ^{+} (v=0) state^{a}’^{b}.The energy levels of the A^{2}Π (v=0) and X^{2}Σ^{+} (v=0) vibronic states were modeled using a traditional effective Hamiltonian approach, where as those for the interacting A^{2}Π (v=1)/B^{2}Σ^{+} (v=0) vibronic levels were modeled by augmenting the traditional effective Hamiltonian with terms to account for local perturbations. A comparison to the analysis of the high temperature spectra performed by Martin^{c}, which did not include many of the spectral features of the present work, will be given.

aC.I. Frum, J.J. Oh, E.A. Cohen, and H.M. Pickett, Ap. J. 408 L61 (1993).

bW.L. Barclay, Jr., M.A. Anderson, and L.M. Zuirys, Ap. J. 408, L65 (1993).

cH. Martin, J. Mol. Spec. 108, 66 (1984).